What is the component concentration ratio, [BrO]/[HBrO], of a buffer that has a pH of 8.02?

(Ka of HBrO = 2.3 × 10−9)
Answer #1

Sol.

As Ka of HBrO = 2.3 × 10-9   

So , pKa = - log(Ka) = - log ( 2.3 × 10-9 ) = 8.64

Using Henderson - Hasselbalch Equation ,

pH = pKa + log ( [BrO-] / [HBrO] )

8.02 = 8.64 + log ( [BrO-] / [HBrO] )  

log ( [BrO-] / [HBrO] ) = 8.02 - 8.64 = - 0.62  

[BrO-] / [HBrO] = 10-0.62 = 0.2399